OpenStudy (anonymous):

The average annual salary of the employees of a company in the year 2005 was $70,000. It increased by the same factor each year and in 2006, the average annual salary was $82,000. Let f(x) represent the average annual salary, in thousand dollars, after x years since 2005. Which of the following best represents the relationship between x and f(x)?

OpenStudy (anonymous):

help mehhhhhhhhh

OpenStudy (anonymous):

@Z4K4R1Y4

OpenStudy (anonymous):

@Z4K4R1Y4 @Z4K4R1Y4 @Z4K4R1Y4

OpenStudy (z4k4r1y4):

where are the options?

OpenStudy (anonymous):

f(x) = 70(1.17)x f(x) = 82(1.17)x f(x) = 70(2.2)x f(x) = 82(2.2)x

OpenStudy (anonymous):

:)

OpenStudy (z4k4r1y4):

lol for the scenario given: x = 1 and f(x) = 82

OpenStudy (z4k4r1y4):

so pick the option where the answer = 82

OpenStudy (z4k4r1y4):

Note: you may have to round up.

OpenStudy (anonymous):

the second one ?

OpenStudy (z4k4r1y4):

for the second one you would have: 82 * 1.17 * 1 = 96

OpenStudy (anonymous):

so it would be the 3rd one

OpenStudy (z4k4r1y4):

the third one would be: 70 * 2.2 * 1 = 154

OpenStudy (anonymous):

sorry 4th on e im stupid

OpenStudy (z4k4r1y4):

show me how you'd do the 4th one

OpenStudy (anonymous):

82 *2.2=17.6 so it would be the second one

OpenStudy (z4k4r1y4):

i get 180

OpenStudy (anonymous):

how._.

OpenStudy (z4k4r1y4):

try the first one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

81.9

OpenStudy (z4k4r1y4):

if you round up what do you get?

OpenStudy (anonymous):

how do i round it up

OpenStudy (anonymous):

is there's like a formula or something

OpenStudy (z4k4r1y4):

if i was to round 59 up to the nearest ten then it would be 60 if iwas to round up 390 to the nearest hundred it would be 400 if i was to round up 6.9 to the nearest whole number it would be 7

OpenStudy (z4k4r1y4):

does that help?

OpenStudy (anonymous):

yesssssss

OpenStudy (z4k4r1y4):

so 81.9 = ?

OpenStudy (anonymous):

82?

OpenStudy (z4k4r1y4):

right!!

OpenStudy (anonymous):

wOoOoOo

OpenStudy (z4k4r1y4):

lol

OpenStudy (anonymous):

haha can you help me with another one .... please pretty please

OpenStudy (z4k4r1y4):

i'll try

OpenStudy (anonymous):

yes! you are awesome!!!!!

OpenStudy (anonymous):

ok here we go The population f(x), in millions, of State A of a country after x years is represented by the function shown below: f(x) = 4(1.08)t The graph shows the population g(x), in millions, of State B of the country after x years: Which conclusion is correct about the population of State A and State B? The original population of State A was double of the original population of State B. The original population of State B was double of the original population of State A. The original population of State A was four times of the original population of State B. The original population of State A was equal to the original population of State B.

OpenStudy (anonymous):

:D

OpenStudy (z4k4r1y4):

by looking at the graph whats the original population of State B?

OpenStudy (anonymous):

um 2?

OpenStudy (z4k4r1y4):

yes. 2 means 2 million.

OpenStudy (anonymous):

ooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooo

OpenStudy (z4k4r1y4):

lol now for state A the equation has the same structure as your previous problem: f(x) = original vaue * constant * number of years

OpenStudy (anonymous):

4.32 :D

OpenStudy (anonymous):

?

OpenStudy (z4k4r1y4):

you can get the original population for state A from the equation without any calculation. there is no need to multiply 4 & 1.08

OpenStudy (z4k4r1y4):



OpenStudy (anonymous):